OpenStudy (anonymous):

find the intercepts of the polynomial x^3+3x^2-x-3

OpenStudy (amoodarya):

x^3+3x^2-x-3= x^3-x+3x^2-3= x(x^2-1)+3(x^2-1) =(x+3)(x^2-1) so x=1 , -1 ,-3

OpenStudy (anonymous):

thanks!